# Interplay between Electroweak Symmetry Breaking and Dark Matter Relic Density

### v2

@@ I Introduction
The spectrum of temperature fluctuations in the Cosmic Microwave Background has offered physicists an extraordinary opportunity to measure the relative abundances of the various energy components of the Universe. The structures that make up these fluctuations are consistent with the existence of a non-baryonic matter component that exceeds the amount of ordinary baryonic matter by a factor around five, offering, together with a number of experimental results from many other experimental sources [1], compelling evidence for the existence of Dark Matter (DM). The best determination of the present DM relic density is [2]
@@
where $\Omega$ is the current energy density in the form of DM relative to the critical energy density of the Universe, and $h$ is today’s reduced Hubble constant in the units of 100100100\,100km/s/MPc.
Amongst the plethora of DM production mechanisms found in the literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], the thermal freeze-out scenario stands out for mirroring the thermal behavior of particles in the Standard Model (SM) [16]. [^1] Namely, photons and neutrinos. This mechanism assumes that in the early stages of our Universe, i.e. at very high temperatures, the interaction rate between DM and particles in the primordial thermal bath was strong enough to keep them in chemical and thermal equilibrium. However, as the Universe expanded and cooled, there arose a moment (the freeze-out instant) when the annihilation rate of DM into lighter particles became subdominant to the Hubble expansion rate, resulting in the decoupling of DM from the thermal plasma and the subsequent freeze-out of its comoving number density.
@@
In practice, analytical and numerical solutions are hard to find, so one usually resorts to approximations to perform the aforementioned integral. One such approximation relies on the assumption that DM particles remain in thermal equilibrium with the primordial bath until the freeze-out instant, implying that the Boltzmann equation becomes dominated by DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale. In effect, this reduces the problem to integrating the Boltzmann Equation only from freeze-out to the present time [17, 18]. Since the relevant integration range lies at lower temperatures, the process becomes effectively low-energy, and finite temperature corrections are typically neglected in this scenario. This contrasts with other DM production mechanisms where these higher-order corrections play an important role in accurately predicting the present DM abundance [19, 20].
@@
The standard approach, therefore, is to calculate all thermally averaged DM depletion processes in the low-energy regime and input them into the Boltzmann Equation. However, this reasoning misses the unavoidable fact that the Higgs boson acquires a non-zero vacuum expectation value (vev), triggering electroweak symmetry breaking (EWSB) at temperatures around $T_{\rm EWSB}\approx 160\,$GeV [21]. It is therefore expected that the particle masses and interaction rates change across the EWSB transition. In this sense, we propose an improved approach for calculating the DM relic density in which both before and after EWSB interactions are considered. As we will demonstrate, neglecting the effects of EWSB in the calculation of the DM relic density is especially concerning for DM candidates that freeze-out before EWSB and can lead to large discrepancies between the standard and improved approaches. We prove that the deviations from the standard approach are driven by a change in the freeze-out instant and/or by a change in the thermally averaged DM depletion processes. In addition, we provide an analytical way to quantify this deviation.
@@
Time-varying masses have been studied before in the context of Variable-Mass Particles, where a scalar DM candidate is coupled to a quintessence field  [22, 23, 24]; in spontaneous freeze-out, where fermionic DM acquires its mass through an early phase transition, and freezes out due to its mass becoming larger than the plasma temperature [25]; in super-cool dark matter, where the universe undergoes a brief period of thermal inflation during which all particles become massless, leading to a dilution of the DM abundance [26]; and in the Flip-Flop vev mechanism, where an auxiliary scalar field acquires a temporary vacuum expectation value, inducing a period where DM is allowed to decay [27]. Furthermore, inclusion of thermally averaged DM depletion processes before and after EWSB has been considered in a model with a light scalar DM candidate produced via freeze-in [28], and in a two-component DM model where one particle freezes in and the other freezes out [29]. Yet, to the best of our knowledge, the effect has not been studied in the case of a single freeze-out DM component, and a systematic comparison between the standard and improved approaches and its physical implications is lacking in the literature.
@@ II.1 Relic density: the standard approach
Consider a SM extension with a single DM candidate of mass $m_{\rm DM}$ which, at high enough temperatures, is kept in thermal and chemical equilibrium with the primordial bath through $2\to 2$ scattering processes. To calculate the present abundance of DM in our Universe, we make use of the Boltzmann Equation to track its number density $n_{\rm DM}$, over time. For practical purposes, it is useful to define the particle number density per comoving volume (or yield) as $Y_{\rm DM}\equiv n_{\rm DM}/s$, where $s$ is the entropy density of the Universe. The Boltzmann equation describing the time evolution of the yield in this scenario can be written as [17, 30]
$\frac{dY_{\rm DM}}{dx}=Z(x)\left(\overline{Y}(x)^{2}-Y_{\rm DM}^{2}\right),$ (2)
where $x\equiv m_{\rm DM}/T$ is the ratio of the DM mass to the temperature of the Universe,
$\overline{Y}(x)=\frac{45}{4\pi^{4}}\frac{x^{2}}{h_{\mathrm{eff}}(x)}K_{2}(x),$ (3)
@@
$Z(x)=\sqrt{\frac{\pi}{45}}\frac{m_{\rm DM}M_{\rm Pl}}{x^{2}}\sqrt{g_{*}}\langle\sigma v\rangle(x).$ (4)
In Eqs. (3) and (4), the function $K_{2}(x)$ is the modified Bessel function of the second kind, $M_{\rm Pl}$ is the Planck mass, $\sqrt{g_{*}}=\frac{h_{\rm eff}}{\sqrt{g_{\rm eff}}}(1+\frac{T}{3h_{\rm eff}}\frac{dh_{\rm eff}}{dT})$, and $h_{\rm eff}$ and $g_{\rm eff}$ are the effective entropy and energy degrees of freedom, respectively. Further, $\langle\sigma v\rangle$ is the total thermal averaged DM annihilation cross-section.
Equation (2) can be solved numerically to find the present DM yield. However, to better grasp the main subject of this paper, it is useful to adopt the semi-analytical approach known as the freeze-out approximation [17, 30], where we assume a freeze-out instant $x_{f}\equiv m_{\rm DM}/T_{f}$, before which DM particles are in thermal equilibrium with the primordial bath (so $Y=\overline{Y}$), and after which they decouple, leading to their comoving number density freezing out (meaning $Y\gg\overline{Y}$). In this sense, it is convenient to reparameterize the yield as $Y_{\rm DM}\approx(1+\delta)\overline{Y}$, where $\delta$ is a monotonically increasing function of $x$. In this approximation, the decoupling instant $x_{f}$ can be numerically obtained by iteration through the following expression [17]
$x_{f}=\log\left(e^{x}\,\frac{\delta_{f}(2+\delta_{f})}{1+\delta_{f}}\frac{Z\,\overline{Y}}{\left(-\frac{\mathrm{d}\overline{Y}}{\mathrm{d}x}\right)}\right)_{x=x_{f}},$ (5)
where $\delta_{f}$ is a $\mathcal{O}(1)$ number [^2] For example, micrOMEGAs [31, 32] uses $\delta_{f}=1.5$. that dictates the amount by which the yield at freeze-out $Y_{f}$ differs from the equilibrium yield at the freeze-out instant $\overline{Y}(x_{f})$, through $Y_{f}=(1+\delta_{f})\overline{Y}(x_{f})$.
After the freeze-out instant, the standard approach is to neglect $\overline{Y}$ in Eq. (2) and integrate it explicitly for any $x>x_{f}$. In turn, the present-day DM yield is given by [17]:
$Y_{\rm DM}^{\rm today}\approx\frac{Y_{f}}{1+Y_{f}A_{x_{f}\to\infty}},$ (6)
where $A_{x_{1}\to x_{2}}$ is a function that quantifies the evolution of the yield between two points $x_{1}$ and $x_{2}$, and is given by
$A_{x_{1}\to x_{2}}=\int_{x_{1}}^{x_{2}}Z(x)\,\mathrm{d}x=\sqrt{\frac{\pi}{45}}m_{\mathrm{DM}}M_{\mathrm{Pl}}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>}{x^{2}}\,\mathrm{d}x.$ (7)
@@ II.2 Relic density: the improved approach
As explained in the previous section, the standard way to calculate the present DM relic density is to integrate the function $Z(x)$ from the $x$ value at decoupling ($x=x_{f}$) until the present epoch ($x\to\infty$). This is consistent with the idea that the Boltzmann equation is dominated by the DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale [18].
However, when EWSB occurs after the DM freeze-out, i.e., when $T_{f}>T_{\rm EWSB}$, special care must be taken in defining Eq. (7). This is because the integrand of Eq. (7) depends on both $m_{\rm DM}$ and $\langle\sigma v\rangle$, whose values change across the EWSB transition.
Note that EWSB is expected to occur at a temperature of approximately $T_{\rm EWSB}\sim 160\,\mathrm{GeV}$ [21], while the freeze-out of a typical scalar DM candidate takes place around $x_{f}\sim 25$ [17]. This implies that EWSB occurs after DM freeze-out when [29]
$x_{f}<x_{\rm EWSB}$$\implies$ $m_{\rm DM}>160\times 25\ {\rm GeV}\sim 4\,{\rm TeV}\,.$ (9)
@@
To properly define the improved approach — which incorporates the effects of EWSB — and to facilitate the comparison with the standard approach — where DM evolution is computed entirely in the electroweak broken phase, as is conventionally done in the dark matter literature —  we must introduce some notation. Let $\sigma\mathrm{v}\right>_{\rm bEWSB}$ be the thermal averaged cross-section before EWSB and $\sigma\mathrm{v}\right>_{\rm aEWSB}$ that after EWSB. Further, let $\widetilde{m}_{\rm DM}$ denote the DM mass in the electroweak unbroken phase and $m_{\rm DM}$ denote the DM mass in the electroweak broken phase. Then, the terminology we will use throughout this paper is
@@
    $\sigma v\right>_{S}$ : represents the total thermal averaged cross-section in the standard approach and is defined by the relation
    $\sigma v\right>_{S}(x)=\left<\sigma v\right>_{\rm aEWSB}(x),$ (10)
    throughout all cosmic history, i.e., over the full range of $x$.
@@
    $\sigma v\right>_{I}$ : denotes the total thermal averaged cross-section in the improved prescription, which differentiates contributions from channels operating before and after EWSB, leading to the relation
[…]
@@
    $Y^{\infty}_{S}=\frac{Y^{S}_{f}}{1+Y^{S}_{f}A^{S}_{{x^{S}_{f}}\to\infty}},$ (12)
    where the DM yield computed at the standard freeze-out instant $x_{f}^{S}$ is
    $Y_{f}^{S}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{S}_{f}\right),$ (13)
@@
    $A^{S}_{{x_{1}}\to x_{2}}=\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\rm DM}\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>_{\rm aEWSB}}{x^{2}}\,\mathrm{d}x.$ (14)
@@
    $Y^{\infty}_{I}=\frac{Y^{I}_{f}}{1+Y^{I}_{f}A^{I}_{{x^{I}_{f}}\to\infty}},$ (15)
    where the DM yield computed at the improved freeze-out instant $x_{f}^{I}$ is
    $Y_{f}^{I}=\left(1+\delta_{f}\right)\overline{Y}\left(x^{I}_{f}\right),$ (16)
@@
    $A^{I}_{{x_{1}}\to x_{2}}$ $=$ $\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,\widetilde{m}_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm bEWSB}}{x^{2}}\,\Theta(x_{\rm{EWSB}}-x)\,\mathrm{d}x$ (17)
    $+\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\,m_{\mathrm{DM}}\,\int_{x_{1}}^{x_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm aEWSB}}{x^{2}}\,\Theta(x-x_{\rm{EWSB}})\,\mathrm{d}x.$

### v3

@@ I Introduction
The spectrum of temperature fluctuations in the Cosmic Microwave Background has offered physicists an extraordinary opportunity to measure the relative abundances of the various energy components of the Universe. The structures that make up these fluctuations are consistent with the existence of a non-baryonic matter component that exceeds the amount of ordinary baryonic matter by a factor around five, offering, together with a number of experimental results from many other experimental sources Cirelli:2024ssz , compelling evidence for the existence of Dark Matter (DM). The best determination of the present DM relic density is Planck:2018vyg
@@
where $\Omega$ is the current energy density in the form of DM relative to the critical energy density of the Universe, and $h$ is today’s reduced Hubble constant in the units of 100100\,km/s/MPc.
Amongst the plethora of DM production mechanisms found in the literature DEramo:2010keq ; Hochberg:2014dra ; Kuflik:2015isi ; DAgnolo:2015ujb ; Pappadopulo:2016pkp ; Garny:2017rxs ; DAgnolo:2020mpt ; Smirnov:2020zwf ; Fitzpatrick:2020vba ; Kramer:2020sbb ; Hryczuk:2021qtz ; Bringmann:2021tjr , the thermal freeze-out scenario stands out for mirroring the thermal behavior of particles in the Standard Model (SM) Bauer:2017qwy . [^1] Namely, photons and neutrinos. This mechanism assumes that in the early stages of our Universe, i.e. at very high temperatures, the interaction rate between DM and particles in the primordial thermal bath was strong enough to keep them in chemical and thermal equilibrium. However, as the Universe expanded and cooled, there arose a moment (the freeze-out instant) when the annihilation rate of DM into lighter particles became subdominant to the Hubble expansion rate, resulting in the decoupling of DM from the thermal plasma and the subsequent freeze-out of its comoving number density.
@@
In practice, analytical and numerical solutions are hard to find, so one usually resorts to approximations to perform the aforementioned integral. One such approximation relies on the assumption that DM particles remain in thermal equilibrium with the primordial bath until the freeze-out instant, implying that the Boltzmann equation becomes dominated by DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale. In effect, this reduces the problem to integrating the Boltzmann Equation only from freeze-out to the present time Cline:2013gha ; Arcadi:2024ukq . Since the relevant integration range lies at lower temperatures, the process becomes effectively low-energy, and finite temperature corrections are typically neglected in this scenario. This contrasts with other DM production mechanisms where these higher-order corrections play an important role in accurately predicting the present DM abundance Baker:2017zwx ; Chakrabarty:2022bcn .
@@
The standard approach, therefore, is to calculate all thermally averaged DM depletion processes in the low-energy regime and input them into the Boltzmann Equation. However, this reasoning misses the unavoidable fact that the Higgs boson acquires a non-zero vacuum expectation value (vev), triggering electroweak symmetry breaking (EWSB) at temperatures around $T_{\rm EWSB}\approx 160\,$GeV Laine:2015kra . It is therefore expected that the particle masses and interaction rates change across the EWSB transition. In this sense, we propose an improved approach for calculating the DM relic density in which both before and after EWSB interactions are considered. As we will demonstrate, neglecting the effects of EWSB in the calculation of the DM relic density is especially concerning for DM candidates that freeze-out before EWSB, i.e. for DM candidates with a mass above $\sim 4$ TeV, and can lead to large discrepancies between the standard and improved approaches. We prove that the deviations from the standard approach are driven by a change in the freeze-out instant and/or by a change in the thermally averaged DM depletion processes. In addition, we provide an analytical way to quantify this deviation.
@@
Time-varying masses have been studied before in the context of Variable-Mass Particles, where a scalar DM candidate is coupled to a quintessence field  Anderson:1997un ; Rosenfeld:2005pw ; Franca:2003zg ; in spontaneous freeze-out, where fermionic DM acquires its mass through an early phase transition, and freezes out due to its mass becoming larger than the plasma temperature Heurtier:2019beu ; in super-cool dark matter, where the universe undergoes a brief period of thermal inflation during which all particles become massless, leading to a dilution of the DM abundance Hambye:2018qjv ; and in the Flip-Flop vev mechanism, where an auxiliary scalar field acquires a temporary vacuum expectation value, inducing a period where DM is allowed to decay Baker:2018vos . Furthermore, inclusion of thermally averaged DM depletion processes before and after EWSB has been considered in a model with a light scalar DM candidate produced via freeze-in Heeba:2018wtf , and in a two-component DM model where one particle freezes in and the other freezes out Bhattacharya:2021rwh . However, to the best of our knowledge, a systematic comparison between the standard and improved approaches and its physical implications is lacking in the literature.
@@ II.1 Relic density: the standard approach
Consider a SM extension with a single DM candidate of mass $m_{\rm DM}$ which, at high enough temperatures, is kept in thermal and chemical equilibrium with the primordial bath through $2\to 2$ scattering processes. To calculate the present abundance of DM in our Universe, we make use of the Boltzmann Equation to track its number density $n_{\rm DM}$, over time. For practical purposes, it is useful to define the particle number density per comoving volume (or yield) as $Y_{\rm DM}\equiv n_{\rm DM}/s$, where $s$ is the entropy density of the Universe. The Boltzmann equation describing the evolution of the yield in this scenario can be written as Cline:2013gha ; Steigman:2012nb
$\frac{dY_{\rm DM}}{dy}=Z(y)\left(\overline{Y}(y)^{2}-Y_{\rm DM}^{2}\right),$ (2)
where [^2] Usually, the independent variable is defined as the ratio of the DM mass to the temperature of the Universe, i.e. $x=m_{\rm DM}/T$. However, since we consider a scenario where $m_{\rm DM}$ changes discontinuously across a phase transition, this definition is ill-defined, and we work with $y=1/T$ instead. $y\equiv 1/T$ is the ratio of the DM mass to the temperature of the Universe,
$\overline{Y}(y)\equiv\frac{45}{4\pi^{4}}\frac{m_{\rm DM}^{2}y^{2}}{h_{\mathrm{eff}}(y)}K_{2}\left(m_{\rm DM}\,y\right),$ (3)
@@
$Z(y)\equiv\sqrt{\frac{\pi}{45}}M_{\rm Pl}\frac{\sqrt{g_{*}}\,\langle\sigma v\rangle}{m_{\rm DM}\,y^{2}}.$ (4)
In Eqs. (3) and (4), the special function $K_{2}$ is the modified Bessel function of the second kind, $M_{\rm Pl}$ is the Planck mass, $\sqrt{g_{*}}=\frac{h_{\rm eff}}{\sqrt{g_{\rm eff}}}(1+\frac{T}{3h_{\rm eff}}\frac{dh_{\rm eff}}{dT})$, and $h_{\rm eff}$ and $g_{\rm eff}$ are the effective entropy and energy degrees of freedom, respectively. Further, $\langle\sigma v\rangle$ is the total thermal-averaged DM annihilation cross-section, computed through the Gondolo-Gelmini formula Gondolo:1990dk :
$\langle\sigma v\rangle\equiv\frac{y}{8m_{\rm DM}^{4}K_{2}^{2}\left(m_{\rm DM}\,y\right)}\int_{4m_{\rm DM}^{2}}^{\infty}\sigma(s-4m_{\rm DM}^{2})\sqrt{s}K_{1}\left(\sqrt{s}\,y\right)\mathrm{d}s.$ (5)
Here, the special function $K_{1}$ is the modified Bessel function of the first kind, and $\sigma$ is the total DM annihilation scattering cross-section.
Equation (2) can be solved numerically to find the present DM yield. All the results and parameter space presented in our study are based on solving the Boltzmann equation numerically, adapted to include the phase transition effects. However, to better grasp the main subject of this paper, it is useful to recall the semi-analytical freeze-out approximation Cline:2013gha ; Steigman:2012nb , which we employ only to demonstrate the subtleties of the analysis. In this approximation, one assumes a freeze-out instant $y_{f}\equiv 1/T_{f}$, before which before which DM particles are in thermal equilibrium with the primordial bath (so $Y=\overline{Y}$), and after which they decouple, leading to their comoving number density freezing out (meaning $Y\gg\overline{Y}$). In this sense, it is convenient to reparameterize the yield as $Y_{\rm DM}\approx(1+\delta)\overline{Y}$, where $\delta$ is a monotonically increasing function of $y$. In this approximation, the decoupling instant $y_{f}$ can be numerically obtained by iteration through the following expression Cline:2013gha
$y_{f}=\log\left(e^{y}\,\frac{\delta_{f}(2+\delta_{f})}{1+\delta_{f}}\frac{Z\,\overline{Y}}{\left(-\frac{\mathrm{d}\overline{Y}}{\mathrm{d}y}\right)}\right)_{y=y_{f}},$ (6)
where $\delta_{f}$ is a $\mathcal{O}(1)$ number [^3] For example, micrOMEGAs Belanger:2001fz ; Alguero:2023zol uses $\delta_{f}=1.5$. that dictates the amount by which the yield at freeze-out $Y_{f}$ differs from the equilibrium yield at the freeze-out instant $\overline{Y}(y_{f})$, through $Y_{f}=(1+\delta_{f})\overline{Y}(y_{f})$.
After the freeze-out instant, the standard approach is to neglect $\overline{Y}$ in Eq. (2) and integrate it explicitly for all $y>y_{f}$. In turn, the present-day DM yield is given by Cline:2013gha :
$Y_{\rm DM}^{\rm today}\approx\frac{Y_{f}}{1+Y_{f}A_{y_{f}\to\infty}},$ (7)
where $A_{y_{1}\to y_{2}}$ is a function that quantifies the evolution of the yield between two points $y_{1}$ and $y_{2}$, and is given by
$A_{y_{1}\to y_{2}}=\int_{y_{1}}^{y_{2}}Z(y)\,\mathrm{d}y=\sqrt{\frac{\pi}{45}}M_{\mathrm{Pl}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>}{m_{\mathrm{DM}}\,y^{2}}\,\mathrm{d}y.$ (8)
@@ II.2 Relic density: the improved approach
As explained in the previous section, the standard way to calculate the present DM relic density is to integrate the function $Z(y)$ from the $y$ value at decoupling ($y=y_{f}$) until the present epoch ($y\to\infty$). This is consistent with the idea that the Boltzmann equation is dominated by the DM depletion processes occurring at temperatures below freeze-out, and hence below the DM mass scale Arcadi:2024ukq .
However, when EWSB occurs after the DM freeze-out, i.e., when $T_{f}>T_{\rm EWSB}$, special care must be taken in defining Eq. (8). This is because the integrand of Eq. (8) depends on both $m_{\rm DM}$ and $\langle\sigma v\rangle$, whose values change across the EWSB transition. Note that EWSB is expected to occur at a temperature of approximately $T_{\rm EWSB}\sim 160\,\mathrm{GeV}$ Laine:2015kra , while the freeze-out of a typical scalar DM candidate takes place around $T_{f}\sim m_{\mathrm{DM}}/25$ Cline:2013gha . This implies that EWSB occurs after DM freeze-out when Bhattacharya:2021rwh
$T_{f}>T_{\rm EWSB}\Leftrightarrow y_{f}<y_{\rm EWSB}$ $\implies$ $m_{\rm DM}>160\times 25\ {\rm GeV}\sim 4\,{\rm TeV}\,.$ (10)
@@
To properly define the improved approach — which incorporates the effects of EWSB — and to facilitate the comparison with the standard approach — where DM evolution is computed entirely in the electroweak broken phase, as is conventionally done in the dark matter literature —  we must introduce some notation. Let $\left<\sigma\mathrm{v}\right>_{\rm bEWSB}$ be the thermal averaged cross-section before EWSB and $\left<\sigma\mathrm{v}\right>_{\rm aEWSB}$ that after EWSB. Further, let $\widetilde{m}_{\rm DM}$ denote the DM mass in the electroweak unbroken phase and $m_{\rm DM}$ denote the DM mass in the electroweak broken phase. Then, the terminology we will use throughout this paper is
@@
    $\left<\sigma v\right>_{S}$ : represents the total thermal averaged cross-section in the standard approach and is defined by the relation
    $\left<\sigma v\right>_{S}(y)=\left<\sigma v\right>_{\rm aEWSB}(y),$ (11)
    throughout all cosmic history, i.e., over the full range of $y$.
@@
    $\left<\sigma v\right>_{I}$ : denotes the total thermal averaged cross-section in the improved prescription, which differentiates contributions from channels operating before and after EWSB, leading to the relation
    $\left<\sigma v\right>_{I}(y)=\left<\sigma v\right>_{\rm bEWSB}(y)\,\Theta(y_{\rm{EWSB}}-y)+\left<\sigma v\right>_{\rm aEWSB}(y)\,\Theta(y-y_{\rm{EWSB}}),$ (12)
@@
    $Y^{\infty}_{S}=\frac{Y^{S}_{f}}{1+Y^{S}_{f}A^{S}_{{y^{S}_{f}}\to\infty}},$ (13)
    where the DM yield computed at the standard freeze-out instant $y_{f}^{S}$ is
    $Y_{f}^{S}=\left(1+\delta_{f}\right)\overline{Y}\left(y^{S}_{f}\right),$ (14)
@@
    $A^{S}_{{y_{1}}\to y_{2}}=\sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{m_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma\mathrm{v}\right>_{\rm aEWSB}}{y^{2}}\,\mathrm{d}y.$ (15)
@@
    $Y^{\infty}_{I}=\frac{Y^{I}_{f}}{1+Y^{I}_{f}A^{I}_{{y^{I}_{f}}\to\infty}},$ (16)
    where the DM yield computed at the improved freeze-out instant $y_{f}^{I}$ is
    $Y_{f}^{I}=\left(1+\delta_{f}\right)\overline{Y}\left(y^{I}_{f}\right),$ (17)
@@
    $A^{I}_{{y_{1}}\to y_{2}}$ $=$ $\sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{\widetilde{m}_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm bEWSB}}{y^{2}}\,\Theta(y_{\rm{EWSB}}-y)\,\mathrm{d}y$ (18)
    $+\ \sqrt{\frac{\pi}{45}}\frac{M_{\mathrm{Pl}}}{m_{\rm DM}}\int_{y_{1}}^{y_{2}}\,\frac{\sqrt{g_{*}}\left<\sigma v\right>_{\rm aEWSB}}{y^{2}}\,\Theta(y-y_{\rm{EWSB}})\,\mathrm{d}y.$